The percentage of the x-21 isotope is 87.99% what is the contribution to the weight average from the x-21 isotope which has a mass of 21.00 amu



1
Expert's answer
2022-02-06T11:19:56-0500

The abundance of X-19 + The abundance of X-21 = 100%

12.01% + The abundance of X-21 = 100%

Hence,

The abundance of X-21 = 100% - 12.01% = 87.99%

The abundance of X-21 = 87.99%


Answer: The percentage of the X-21 isotope is 87.99%.